# FuncGenFoil: Airfoil Generation and Editing Model in Function Space

### v1

@@ Abstract
Aircraft manufacturing is the jewel in the crown of industry, among which generating high-fidelity airfoil geometries with controllable and editable representations remains a fundamental challenge. While existing deep-learning-based methods rely on predefined parametric function families, e.g., Bézier curves and discrete point-based representations, they suffer from inherent trade-offs between expressiveness and resolution flexibility. To tackle this challenge, we introduce FuncGenFoil, a novel function-space generative model that directly learns functional airfoil geometries. Our method inherits both the advantages of arbitrary resolution sampling and the smoothness of parametric functions, as well as the strong expressiveness of discrete point-based functions. Empirical evaluations on the AFBench dataset demonstrate that FuncGenFoil improves upon state-of-the-art methods in airfoil generation by achieving a relative -74.4% label error reduction and +23.2% diversity increase on the AF-200K dataset. Our results highlight the advantages of function-space modeling for aerodynamic shape optimization, offering a powerful and flexible framework for high-fidelity airfoil design. Our code will be released.
@@ 1 Introduction
The airfoil inverse design problem serves as a central aspect of aircraft manufacturing. Traditionally, given geometric requirements, engineers first select the most similar airfoils from well-known airfoil datasets (e.g., NACA Cummings et al. (2015)) and leverage the trial-and-error strategy Sharma et al. (2021). Considering the mission of the aircraft, an initial airfoil design that meets the design conditions is preliminarily created. Then, through rounds of physical analysis, such as aerodynamics and mechanics, the airfoil is iteratively optimized to achieve better performance until the requirements are met. In practice, such direct design procedures are highly inefficient and time-consuming, often taking months. To minimize development and design time, as well as associated costs, automatic design methods have been introduced as efficient alternatives in aircraft manufacturing engineering. In particular, machine learning-based design and optimization techniques have gained significant attention. However, before applying algorithms to airfoil design, it is crucial to determine appropriate methods for representing airfoils within these algorithms.
Existing methods for airfoil representation can generally be divided into two categories: parametric-model-based approaches Xie et al. (2024) and discrete-point-based methods Liu et al. (2024); Sekar et al. (2019). First, parametric-model-based methods predefine function families, e.g., Bézier curves Chen and Fuge (2021), Hicks-Henne curves, and NURBS, and leverage mathematical optimization techniques or generative models to determine the parameters of these functions for a new airfoil. These methods rigorously preserve key geometric properties of the defined functional families, e.g., high-order smoothness. Furthermore, such functional representations of airfoils allow for arbitrary sampling of control points in real manufacturing, given the constraints of engineering precision. Despite these benefits, parametric-model-based methods suffer from a significantly reduced design space, i.e., selecting a specific function family excludes the possibility of other shapes, which limits the upper bound of airfoil design algorithms. Second, discrete-point-based methods directly generate multiple points to represent airfoil shapes. These methods maximize the design space of airfoils but cannot maintain some important mathematical properties, e.g., continuity. Furthermore, they cannot directly generate control points at arbitrary resolutions because the number of generated points is typically fixed for each model after training.
@@
In this paper, we answer this question by proposing FuncGenFoil, a novel function-space generative model for airfoil representation (see Fig. 1). Different from previous data-driven and generative methods, e.g., cVAE Kingma (2013) and cGAN Mirza (2014), which directly generate discrete points, our approach models airfoil geometry using a general continuous function approximator Anandkumar et al. (2019); Li et al. (2021); Kovachki et al. (2023); Azizzadenesheli et al. (2024), i.e., neural operator architectures. Simultaneously, it leverages recent advances in generative methods, e.g., diffusion models Ho et al. (2020a) and flow matching models Lipman et al. (2022), to generate diverse airfoils beyond the design space of any predefined geometry function. Our method has the advantages of both parametric-model-based approaches and discrete-point-based methods. Due to its functional representation, the generated airfoil is continuous and can be sampled at arbitrary resolutions, making it easier to manufacture. Moreover, thanks to the general functional approximator nature of neural operators, our method can explore a broader design space beyond predefined functional families.
Specifically, we leverage the flow matching framework Lipman et al. (2022), an improved alternative to diffusion models, and FNO Li et al. (2021), a resolution-free neural operator, as the backbone of our generative model to design FuncGenFoil. In the forward process, we perturb the airfoil into a noise distribution through straight flows. In the backward process, we reconstruct the airfoil by reversing the flow direction. By minimizing the distance between the reconstructed airfoils and actual airfoils, we learn the neural operators that enable us to generate airfoils from functions of Gaussian process. Beyond generation, our method also supports airfoil editing by incorporating the airfoils to be edited as conditions for the generative model.
In summary, our main contributions are threefold: (1) We propose generating airfoil shapes in the functional space to achieve important properties for the aircraft engineering, i.e., arbitrary-resolution control point sampling and maximal design space; (2) We design FuncGenFoil, the first controllable airfoil generative model in the functional space, which effectively incorporates neural operator architectures into the generative model; (3) We further enhance FuncGenFoil with airfoil editing capabilities through minimal adaptations. Experimental results on the AFBench dataset indicate that our proposed method achieves state-of-the-art airfoil generation quality, reducing label error by 74.4% and increasing diversity by 23.2% on the AF-200K dataset, as validated by aerodynamic simulation analysis. In addition, our method is the first to successfully perform airfoil editing by fixing and dragging at any position, achieving nearly zero MSE error (less than $10^{-7}$).
@@ 2 Related Works
Generative Models. Generative models based on score matching Ho et al. (2020b); Song et al. (2021) and flow matching Lipman et al. (2023); Tong et al. (2024) have significantly advanced machine learning, achieving state-of-the-art results in areas such as image generation Rombach et al. (2022), text generation Gat et al. (2024), and video generation Ho et al. (2022); Polyak et al. (2024). However, most of these models operate in finite-dimensional spaces and rely on fixed discretizations of data. Such formulations hinder transferability across discretizations and ignore function-level constraints (e.g., continuity, smoothness), motivating the need for function-space generative modeling.
@@
Neural Architecture for Functions space Designing neural architectures to handle function spaces remains a major research challenge. Standard networks typically assume fixed-size inputs, making them unsuitable for arbitrary resolutions. Implicit Neural Representations like SIREN Sitzmann et al. (2020) harness random Fourier features Rahimi and Recht (2007) to represent continuous and differentiable objects through position embeddings. Similarly, NeRF Mildenhall et al. (2021) treats input coordinates as continuous variables, offering flexible resolution for function outputs. Neural operators Anandkumar et al. (2019); Li et al. (2021); Kovachki et al. (2023); Azizzadenesheli et al. (2024) and Galerkin transformers Cao (2021) further generalize neural architectures to process sets of points as functional inputs, enabling function-space learning.
@@
Generative Models in Function Space. In generative modeling, early Neural Processes Garnelo et al. (2018b, a) drew on Gaussian Processes Rasmussen (2003), and later methods such as GASP Dupont et al. (2022b), Functa Dupont et al. (2022a), and GANO Rahman et al. (2022) treat data as function evaluations to enable discretization-independent learning. Energy-based and diffusion models Lim et al. (2023); Kerrigan et al. (2023); Lim et al. (2025); Pidstrigach et al. (2023); Franzese et al. (2024) along with flow-based approaches like FFM Kerrigan et al. (2024) and OpFlow Shi et al. (2024) further extend these ideas. Ultimately, creating comprehensive generative models in function space requires defining suitable stochastic processes, score operators, and consistent neural mappings with specialized training for numerical stability—challenges that remain largely unresolved.
Figure 2: Overview of FuncGenFoil’s neural network. The model is a Fourier Neural Operator, but any other neural operator capable of general-purpose function-space approximation may be used. The model takes as input a function $u_{t}$ (at any resolution $d$), design condition variables $c$, and the generation time $t$. It then processes these as an operator and outputs the current velocity operator $v_{\theta}(u_{t},c,t)$ of the ODE for generation.
Airfoil Design and Optimization. Airfoil design is critical for aerodynamic performance in systems such as aircraft, race cars, and wind turbines. Geometric parameterization enables efficient modeling and optimization of airfoil shapes. Free-Form Deformation (FFD) Mortenson (1999), B-splines Farin (2002), or NURBS Schoenberg (1964)—and is used in CAD/CAE tools like ESP Dannenhoffer (2024) and OpenVSP Hahn (2010)—but its independent control points can lead to local optima, instability, and high dimensionality. Modal parameterization methods, including Proper Orthogonal Decomposition Berkooz et al. (1993), global modal Brüls et al. (2007), and compact modal approaches Li and Zhang (2021), address dimensionality by capturing global features with fewer variables, yet their linear reduction struggles with large deformations and detailed edits. In contrast, the Class-Shape Transformation (CST) Kulfan (2008) offers intuitive physical interpretation and differentiability but is limited in handling significant deformations and is sensitive to parameter selection. Thus, a high-degree-of-freedom geometric representation is needed to support arbitrary shape modifications while accurately capturing highly nonlinear deformations with stability and robustness.
Recently, with the rise of generative models in the AI field, airfoil design has increasingly adopted generative models such as VAE, GAN and diffusion models Chen and Fuge (2021); Li et al. (2022); Xie et al. (2024); Yang et al. (2024); Wei et al. (2024a); Liu et al. (2024), as well as CFD meshes Wei et al. (2023, 2024b), to enhance the representation of airfoil shapes in a generative manner. These models establish an effective transformation from a latent space to the airfoil space. Previous works have primarily focused on modeling the discrete points of airfoils, which limits their flexibility for downstream tasks.
@@ 3 FuncGenFoil: Function-space Generative Model for Airfoils
Different from existing airfoil generation methods, FuncGenFoil is constructed as a function-space generative model, capable of producing airfoil geometries as continuous functions instead of discrete points, leveraging the advantages of both parametric-model-based and discrete-point-based methods. We detail the process for airfoil generation and editing tasks, respectively.
@@ 3.1 Airfoil Generation
The entire FuncGenFoil is essentially an ordinary differential equation (ODE) generative model that generates airfoil by solving an ODE continuous time $t\in[0,1]$ as follows:
$\mathrm{d}u_{t}=v_{t}\,\mathrm{d}t,$ (1)
where $v_{t}$ is a velocity that gradually transforms an assumed latent coding $u_{0}$ ($u_{t}$ at $t=0$) sampled from stochastic process $\mathcal{P}$ into an airfoil function $u_{1}$ belonging to the target airfoil distribution $\mathcal{Q}$.
Velocity Operator. Obviously, the velocity $v_{t}$ is the key, we achieve it by establishing a parameterized velocity operator $v_{\theta}(u_{t},c,t)$ with model weights $\theta$, i.e., $v_{t}=v_{\theta}(u_{t},c,t)$. To enable the velocity operator $v_{\theta}$ an ability to take the airfoil shape function as input, we build $v_{\theta}$ as a function-space generative model by Operator Flow Matching Shi et al. (2025) as shown in Fig. 2. Specifically, $v_{\theta}$ consists of stacked multiple Fourier Neural Operator layers and it takes noised airfoil $u_{t}$ (at any resolution $d$), condition variables (optional) and a timestamp $t$ as inputs, deriving $v_{t}$.
Figure 3: Airfoil generation by FuncGenFoil. We first sample latent function from a zero-mean Gaussian process with kernel function $K$ at any resolution. Then we solve an ODE flow through $t=0$ to $t=1$ using numerical ODE solver using velocity operator $v_{\theta}$ along this generation process and can generate diverse airfoils. As shown in (c), A single FuncGenFoil model can generate airfoils at arbitrary resolutions, from low to high, while guaranteeing identical outputs when sampling from the same latent function—an inherent property of the model’s architecture.
Airfoil Geometry Parametrization. Since an airfoil has circular topology, we introduce the variable $\alpha\sim[0,2\pi]$ for convenience. We denote $y(\alpha)=u_{1}(\alpha)$ and $x(\alpha)=\frac{\cos(2\pi\alpha)+1}{2}$ [^1] For the remainder of this paper, we omit the explicit dependence on $\alpha$ for clarity., achieving the function of airfoil $u_{0}=(x(\alpha),y(\alpha))$.
@@
Training. We train $v_{\theta}$ under the denoising training process. Given an airfoil geometry $u_{1}$, we compute its corresponding noised sample at time $t$ as follows:
$u_{t}=t\times u_{1}+(1-t)\times u_{0}$ (2)
@@
Then we feed $u_{t}$ into $v_{\theta}$ and train $v_{\theta}$ by matching the velocity operator:
$\mathcal{L}_{\mathrm{FM}}=E_{t\sim[0,1],u_{0},u_{1}}\left[\|v_{t}-v_{\theta}(u_{t},c,t)\|^{2}\right],$ (3)
where $v_{t}$ is the ground-truth velocity:
$v_{t}=u_{1}-u_{0}.$ (4)
Detailed algorithm for training velocity operator is shown in Appendix B Algorithm 2.
@@
Inference. Given a trained velocity operator $v_{\theta}$, the inference process, i.e., airfoil generation process, is equal to deriving airfoil geometry at time $t=1$, denoted as $u_{1}$, based on a latent coding $u_{0}$ sampled from the stochastic process $\mathcal{P}$. $\mathcal{P}$ is assumed as a Gaussian Process $\mathcal{GP}(0,K)$, where $K$ is a covariance kernel function. Therefore, $u_{1}$ could be derived by solving the ODE in Equation 1 numerically (e.g., using the Euler method) as follows:
@@
The generation process is shown in Fig. 3 while the detailed algorithm is shown in Appendix B Algorithm 3.

### v3

@@ Abstract
Aircraft manufacturing is the jewel in the crown of industry, in which generating high-fidelity airfoil geometries with controllable and editable representations remains a fundamental challenge. Existing deep learning methods, which typically rely on predefined parametric representations (e.g., Bézier curves) or discrete point sets, face an inherent trade-off between expressive power and resolution adaptability. To tackle this challenge, we introduce FuncGenFoil, a novel function-space generative model that directly reconstructs airfoil geometries as function curves. Our method inherits the advantages of arbitrary-resolution sampling and smoothness from parametric functions, as well as the strong expressiveness of discrete point-based representations. Empirical evaluations demonstrate that FuncGenFoil improves upon state-of-the-art methods in airfoil generation, achieving a relative 74.4% reduction in label error and a 23.2% increase in diversity on the AF-200K dataset. Our results highlight the advantages of function-space modeling for aerodynamic shape optimization, offering a powerful and flexible framework for high-fidelity airfoil design.
@@ 1 Introduction
The airfoil inverse design problem serves as a central aspect of aircraft manufacturing. Traditionally, given geometric requirements, engineers first select the most similar airfoils from well-known airfoil datasets (e.g., NACA [8]) and rely on a trial-and-error strategy [51]. Considering the mission of the aircraft, an initial airfoil design that meets these conditions is preliminarily created. Then, through iterative rounds of physical analyses, such as aerodynamic and mechanical evaluations, the airfoil is optimized to achieve improved performance until the specified requirements are met. In practice, such direct design procedures are highly inefficient and time-consuming, often taking months. To minimize development and design time, as well as associated costs, automated design methods have been introduced as efficient alternatives in aircraft manufacturing engineering. In particular, machine learning-based design and optimization techniques have gained significant attention. However, before applying these algorithms to airfoil design, it is crucial to determine appropriate methods for representing airfoils within these algorithms.
@@
Existing methods for airfoil representation can generally be divided into two categories: parametric-model-based approaches [61] and discrete-point-based methods [37, 49]. First, parametric-model-based methods predefine function families, e.g., Bézier curves [7], Hicks-Henne curves, and NURBS, and leverage mathematical optimization techniques or generative models to determine the parameters of these functions for generating a new airfoil. These methods rigorously preserve key geometric properties of the defined function families, e.g., higher-order smoothness. Furthermore, such functional representations of airfoils allow arbitrary sampling of control points in real manufacturing, given engineering precision constraints. Despite these benefits, parametric-model-based methods suffer from a significantly reduced design space; i.e., selecting a specific function family excludes other possible shapes, thus limiting the upper bound of airfoil design algorithms. Second, discrete-point-based methods directly generate multiple points to represent airfoil shapes. These methods maximize the airfoil design space but cannot inherently maintain certain important mathematical properties, e.g., continuity. Furthermore, they cannot directly generate control points at arbitrary resolutions, as the number of generated points is typically fixed for each model after training.
@@
In this paper, we address this question by proposing FuncGenFoil, a novel function-space generative model for airfoil representation (see Figure 1). Unlike previous data-driven generative methods in finite-dimensional spaces, e.g., cVAE [25], cGAN [41], diffusion models [19], or flow matching models [35], which directly generate discrete point outputs, we leverage recent advances in diffusion and flow models defined in function space [33, 23, 32, 42, 13, 52] to produce infinite-dimensional outputs as general continuous functions. Simultaneously, our approach models airfoil geometry using Neural Operator architectures [1, 31, 27, 2], enabling the generation of diverse airfoils beyond the design spaces of predefined geometric function families, thanks to its general operator approximation capability. Our method combines the advantages of both parametric-model-based and discrete-point-based approaches. Due to its intrinsic representation in function space, the generated airfoils are continuous, and both training and sampling can be performed at arbitrary resolutions, facilitating downstream optimization and manufacturing processes.
Specifically, we use the flow matching framework [35], an improved alternative to diffusion models, and FNO [31], a resolution-free neural operator, as the backbone of our generative model to design FuncGenFoil. During training, we perturb the airfoil into a noise function through straight flows and let neural operators learn the airfoil’s deconstructed velocity via flow matching. During inference, we reconstruct the airfoil by reversing the flow direction starting from a Gaussian process. Beyond generation, our method also supports airfoil editing, generating new airfoils from an original airfoil by conforming to certain geometric constraints. This can be accomplished by a few steps of fine-tuning with no additional data other than the original airfoil. Constraints can be provided in various forms, such as specifying contour points that the airfoil must pass through or setting geometric constraint parameters, such as thickness.
In summary, our main contributions are threefold: (1) We propose generating airfoil shapes in function space to achieve important properties for aircraft engineering, i.e., arbitrary-resolution control point sampling and maximal design space; (2) We design FuncGenFoil, the first controllable airfoil generative model in function space, which effectively incorporates neural operator architectures into the generative modeling framework; and (3) We further enhance FuncGenFoil with airfoil editing capabilities through minimal adaptations. Experimental results indicate that our proposed method achieves state-of-the-art airfoil generation quality, reducing label error by 74.4% and increasing diversity by 23.2% on the AF-200K dataset, as validated through aerodynamic simulation analysis. In addition, our method is the first to successfully perform airfoil editing by fixing or dragging points at arbitrary positions, achieving nearly zero MSE (less than $10^{-7}$).
@@ 2 Related Works
Generative Models. Generative models based on score matching [20, 55] and flow matching [34, 57, 36] have significantly advanced machine learning, achieving state-of-the-art results in areas such as image generation [47], text generation [16], and video generation [21, 43]. However, most of these models operate in finite-dimensional spaces and rely on fixed discretizations of the data. Such formulations hinder transferability across discretizations and neglect function-level constraints (e.g., continuity, smoothness), motivating the need for generative modeling in function space.
@@
Neural Architectures for Function Modeling. Designing neural architectures to handle function spaces remains a major research challenge. Standard networks typically assume fixed-size inputs, making them unsuitable for arbitrary resolutions. Implicit neural representations, such as SIREN [54], harness random Fourier features [44] to represent continuous and differentiable objects through position embeddings. Similarly, NeRF [40] treats input coordinates as continuous variables, offering flexible resolution for function outputs. Neural operators [1, 31, 27, 2] and Galerkin transformers [5] further generalize neural architectures to process sets of points as functional inputs, enabling function-space learning. Recently, a neural operator with localized integral and differential kernels has been introduced to improve its capability in capturing local features [38]. To handle complex and variable geometries, the point cloud neural operator has been proposed, which adaptively processes point cloud datasets [63].
@@
Generative Models in Function Space. Early Neural Processes [15, 14] drew upon Gaussian processes [46], and later methods such as GASP [11], Functa [10], and GANO [45] treat data as function evaluations to enable discretization-independent learning. Energy-based and diffusion models [33, 23, 32, 42, 13], along with flow-based approaches like FFM [24] and OpFlow [52], further extend these ideas. Ultimately, developing comprehensive generative models in function space requires defining suitable stochastic processes, score operators, and consistent neural mappings, along with specialized training methods for numerical stability—challenges that remain largely unresolved.
Airfoil Design and Optimization. Airfoil design is essential for aircraft and wind turbines. Geometric parameterization supports efficient shape modeling and optimization. Techniques such as Free-Form Deformation (FFD) [12] and NURBS [48] are widely used in CAD/CAE tools [9, 18], but their independent control points can cause instability, high dimensionality, and suboptimal solutions. Modal methods, including Proper Orthogonal Decomposition [3], global modal [4], and compact modal [30] approaches, reduce dimensionality by encoding global features, though they struggle with large or detailed shape changes. Class-Shape Transformation (CST) [28] offers interpretable and differentiable shapes but lacks flexibility for large deformations and is sensitive to parameter choices. Therefore, a high-degree-of-freedom representation is needed for robust, nonlinear shape modeling. Unlike general computer vision tasks [56, 26], engineering AI models are domain-specific and often limited by sparse data. Generative models are advantageous here, as they can exploit unlabeled or untested samples. In airfoil design, VAEs, GANs, and diffusion models [7, 29, 61, 62, 58, 37], along with CFD-based mesh representations [59, 60], have been used to map latent spaces to airfoil shapes. However, most existing methods model discrete airfoil points, limiting flexibility for downstream applications.
@@ 3 FuncGenFoil: Function-Space Generative Model for Airfoils
In contrast to existing airfoil generation methods, FuncGenFoil is constructed as a function-space generative model capable of producing airfoil geometries as continuous functions rather than discrete points, thereby leveraging the advantages of both parametric-model-based and discrete-point-based methods. In this section, we detail the processes for airfoil generation and editing tasks, respectively.
@@ 3.1 Airfoil Generation
Airfoil Parametrization. Since an airfoil curve $(x,y)$ has circular topology, we introduce the variable $\alpha\in[0,1]$ as the domain of the function, denoting $y(\alpha)=f(x(\alpha))$ and $x(\alpha)=\frac{\cos(2\pi\alpha)+1}{2}$, [^1] For the remainder of this paper, we omit the explicit dependence on $\alpha$ for clarity and convenience. as shown in Figure 7. The entire FuncGenFoil framework is essentially an ordinary differential equation (ODE)-based generative model that produces airfoil curve functions by solving an ODE along continuous time $t\in[0,1]$ as $u_{1}=u_{0}+\int_{t=0}^{t=1}v_{t}\,\mathrm{d}t$, where $u_{1}=y$, starting from some latent function $u_{0}$.
Velocity Operator. The velocity operator $v_{t}$ gradually transforms a latent function $u_{0}$ sampled from a stochastic process $\mathcal{P}$ into an airfoil function $u_{1}$ belonging to the target airfoil distribution $\mathcal{Q}$. The velocity operator is a key component, as it must handle function inputs consisting of point sets with arbitrary resolution or positions. Specifically, the operator takes as input the partially generated airfoil $u_{t}=[u_{t}(\alpha_{0}),u_{t}(\alpha_{1}),u_{t}(\alpha_{2}),\dots]$ and outputs a velocity function corresponding to these positions, $[v_{t}(\alpha_{0}),v_{t}(\alpha_{1}),v_{t}(\alpha_{2}),\dots]$. We realize this by establishing a parameterized neural operator $v_{\theta}$ with model weights $\theta$, constructed as a Neural Operator model capable of processing function-space data at arbitrary resolutions in a single unified model, as shown in Figure 2. Specifically, FuncGenFoil acts as a conditional continuous-time generative model, where the velocity operator $v_{\theta}(u_{t},c,t)$ takes the noised airfoil $u_{t}$, optional conditioning variables $c$, and generation timestamp $t$ as inputs, and consistently outputs the deformation $v_{t}$ as a function. We train $v_{\theta}(u_{t},c,t)$ using Operator Flow Matching [53].
Figure 2: Top: Overview of FuncGenFoil’s neural network and training scheme. The model is a Fourier Neural Operator designed for point cloud data, although any other neural operator capable of general-purpose function-space approximation may be used. The model takes as input a function $u_{t}$ (point cloud data at an arbitrary resolution $d$), optional design condition variables $c$, and the generation time $t$. It then processes this input function consistently and outputs the current velocity operator $v_{\theta}(u_{t},c,t)$ for calculating the flow matching loss. Bottom: Inference with FuncGenFoil is conducted by first sampling a random latent function from a Gaussian process and then reconstructing the airfoil by solving an ODE.
@@
Training. We train $v_{\theta}$ under the simplest denoising training process. Given an airfoil geometry $u_{1}$, we compute its corresponding noised sample at time $t$ as follows:
$u_{t}|_{u_{0},u_{1}}=t\times u_{1}+(1-t)\times u_{0}$ (1)
and then the ground-truth velocity $v_{t}(u_{t})|_{u_{0},u_{1}}$ given $u_{0}$ and $u_{1}$ can be computed as:
$v_{t}(u_{t})|_{u_{0},u_{1}}=\frac{du_{t}}{dt}=u_{1}-u_{0}.$ (2)
@@
We can train $v_{\theta}$ by matching the velocity operator using Flow Matching loss:
$\mathcal{L}_{\mathrm{FM}}=E_{t\sim[0,1],u_{t}}\left[\|v_{t}(u_{t})-v_{\theta}(u_{t},c,t)\|^{2}\right],$ (3)
which according to Conditional Matching Theorem, has same gradients with Conditional Flow Matching loss:
$\mathcal{L}_{\mathrm{CFM}}=E_{t\sim[0,1],u_{0},u_{1}}\left[\|v_{t}(u_{t})|_{u_{0},u_{1}}-v_{\theta}(u_{t}|_{u_{0},u_{1}},c,t)\|^{2}\right],$ (4)
$\nabla_{\theta}\mathcal{L}_{\mathrm{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\mathrm{CFM}}(\theta).$ (5)
See more proofs in Chapter 4 of Flow Matching Guide and Code [36].
@@
Inference. Given a trained velocity operator $v_{\theta}$, the inference process, i.e., airfoil generation process, is equal to deriving airfoil geometry at time $t=1$, denoted as $u_{1}$, based on a latent coding $u_{0}$ sampled from the stochastic process $\mathcal{P}$. $\mathcal{P}$ is assumed as a Gaussian Process $\mathcal{GP}(0,K)$ in this work, where $K$ is a covariance kernel function. Therefore, $u_{1}$ could be derived by solving the generation ODE numerically (e.g., using the Euler method) as follows:
@@
Detailed implementations for training velocity operator and model inference are shown in Appendix B, Algorithm 1 and Algorithm 2.